a telegraph pole is placed on a two wheeled folly located 4m from the thicker end and an upward force of 1500 N at the thinner end is required to keep it horizontal. the pole is 15m long and weighs 9000n, where is the centre of gravity
1
Expert's answer
2018-02-09T08:54:08-0500
Fup = Fdown
Fthin end + Ftruck = Fpole
1500 N + F = 9000 N
Ftruck = 7500 N
The torques at the truck and at thin end isT1, while the torque at center of gravity is T2.
T1 = T2
T = F d
7500 N x (4 m) + 1500 x (15 m) = 9000 N x d
9000 d= 30000 + 22500
9000 d= 52500
d = 52500 / 9000
d = 5.8 m